OpenStudy (anonymous):

Dividing rational expressions. please solve and simplify n^2-n-6 over/ 14 divided by n-3

OpenStudy (anonymous):

Can you use the equation button to give it in a better form? As you've presented it that equation is fairly ambiguous.

OpenStudy (anonymous):

@Indivicivet i'm not sure how, i will try now

OpenStudy (anonymous):

\[\frac{ n ^{2}-n-6}{ 14 } \div n-3\]

OpenStudy (anonymous):

@Indivicivet I did :)

OpenStudy (anonymous):

And that n - 3 is not supposed to be in brackets?

OpenStudy (anonymous):

@Indivicivet this is how it is written in my book. I think you're supposed to make it \[\frac{ n-3 }{ 1 }\]

OpenStudy (anonymous):

Well, that would just be putting it in brackets. In that case it would be: \[\frac{n^2-n-6}{14} \div (n-3) = \frac{(n-3)(n+2)}{14(n-3)} = \frac{n+2}{14}\] I think it's safe to assume that's how the question is supposed to be interpreted.

OpenStudy (anonymous):

@Indivicivet thank you! I see where I was making my mistake now.